# Practical X-ray Gastric Cancer Screening Using Refined Stochastic Data Augmentation and Hard Boundary Box Training

### v3

@@ 2.3 Hard negative mining
Hard sample or hard example mining, which improves model performance by explicitly retraining the model on error-prone or hard-to-identify data (hard samples), is widely used in machine learning (Felzenszwalb et al., 2009; Shrivastava et al., 2016). It is an effective method for counteracting the imbalance of each class and has been widely used, together with metric learning, for face or person detection from complex backgrounds (Zhang et al., 2016; Smirnov et al., 2018; Chen et al., 2020) and other applications.
In recent years, deep learning-based object detection models have also incorporated this strategy and reported many performance improvements by relearning misdetected bounding box regions, with results in the medical field (Li et al., 2019; Tang et al., 2019). However, such object detection models cannot use images that do not contain a single class of detection targets (such as background images or medical healthy control images) for training. This is because the error function to be minimized is defined based on the position and size of the bounding box of each detection target. Existing hard sample mining methods have been applied to data with at least one annotation box of the class to be detected and have used overdetected or undetected regions or image patches as hard samples. Our HBBT is a hard sample mining or hard negative mining; that is, it retrains regions of false positives but differs significantly in that it allows the use of healthy control cases, which could not be used to train object detection models, for training. This is a particularly important contribution in the medical field, where the cost of positive samples is high.
@@ 3 Practical gastric cancer screening system
This study proposed a practical gastric cancer screening system using gastric X-ray images. The system displayed suspected cancerous sites with bounding boxes so physicians can confirm the validity of the estimation. The two technical proposals in this study significantly improved the performance of the system and achieved practical accuracy. The overall structure of the proposed system is shown in Figure 1. This study was approved by the Institutional Review Board of Tokai University Hospital (No. 20R-033, date of approval: June 12, 2020). The main technical proposals were (1) R-sGAIA to effectively detect gastric fold regions and (2) HBBT to avoid false positives by assigning “hard-sample” classes to the bounding box regions where false positives are detected and iteratively learning them. The object detection model used is EfficientDet (Tan et al., 2020), which uses EfficientNet (Tan and Le, 2019), a CNN with reported excellent performance, as its backbone, but the choice of the model used was arbitrary. The input was 2,048 $\times$ 2,048 pixels, and the output was a bounding box surrounding the presumed cancerous region in the image and its confidence score.
@@ 3.1 R-sGAIA
The proposed R-sGAIA is a refined version of our previous proposal, sGAIA (Okamoto et al., 2019). The proposed R-sGAIA and sGAIA are probabilistic online data augmentation methods that can generate various images highlighting gastric folds in X-ray images based on medical knowledge to detect inflammation on the gastric mucosal surface. The former sGAIA improved the detection of gastric cancer by 6.9% in terms of the F1-score (recall 92.3%, precision 32.4%) on a Faster R-CNN model with ResNet-101 as the backbone, validating its practical screening performance. However, in the sGAIA, the emphasis probability of each pixel $p(x,y)$, which is important for an effective enhancement, was subjectively determined as a discrete value based on the results of preliminary experiments. In the proposed R-sGAIA, $p(x,y)$ becomes a flexible continuous function based on only two hyperparameters.
A schematic of the proposed R-sGAIA process is shown in Figure 2. Like sGAIA, R-sGAIA consists of the following four steps, where only step 2 is different.
@@
    $=$ $\frac{1}{1+\exp(-\gamma(E(x,y)+\theta))}.$
@@
Here, $\alpha$ and $\beta$ are augmented parameters. The value of $\alpha$ ranges from 0.9 to 1.0 and that of $\beta$ ranges from --15 to --5, and they are chosen randomly for each processing. Based on this, the proposed R-sGAIA probabilistically enhances the location and intensity of the gastric fold region each time. Thus, it produces a different enhanced gastric image for each trial and is highly compatible with the online data augmentation in deep neural network training. Another significant advantage of the proposed R-sGAIA is that it does not require any additional runtime. We used this R-sGAIA for online augmentation of a machine learning model (i.e., EfficientDet-D7) used for cancer region detection.
@@ 3.2 HBBT
The diagnostic support system for gastric X-ray images is a one-class detection problem that targets only positive (i.e., malignant) regions. However, our task has many regions with similar image features, making them difficult to discriminate even for experts. Given a large number of training images, the recently developed deep learning-based object detection algorithms can achieve the desired detection performance for the task, but this is difficult in situations where the labeled supervisory data is limited. The images in Figure 4 are typical false positives indicated by experts in experiments conducted using a vanilla detection model (i.e., EfficientDet-D7 without the proposed HBBT). The results show that most of these false positives fall broadly into the five categories shown in examples (b)–(f) in Figure 4. However, it is difficult to clearly define and control these features in advance. Specifically, in a fine-grained task like our current goal, it is very important to reduce such hard negatives that have high image similarity but are not the target.
@@
One promising countermeasure to this problem is hard sample mining, of which hard negative mining, which aims to reduce overdetection, is consistent with this goal and has been extensively studied as mentioned above. However, the training of hard negative mining applied to object detection models is aimed at reducing false positives (hard negatives) outside the target region in the training image, and the data that can be used for training are subject to the constraints of the object detection model. In other words, the training data are only images with some object detection annotation label, unless a framework such as semi-supervised learning is used.
@@
In fact, as mentioned earlier, the biggest room for improvement in object detection models is the active use of negative samples, i.e., healthy controls. Conventional object detection algorithms based on deep learning only use images in which there is an object to be detected, and control cases without an object to be detected cannot be included in the training. This is a very important limitation. Therefore, for more effective training of object detection models, we propose HBBT, a very simple and effective method for suppressing false positives, referring to the concept of hard negative mining. In general, training negative samples, which are more prone to false positives among the training data, is effective in constructing detectors with few false positives. In HBBT, as shown in Figure 1, the false-positive boundary boxes detected by the model are called hard boundary boxes and are given hard-sample labels to train the model as a two-class detection problem. This newly added hard-sample class label is treated as a negative example. Training is repeated until an optimal F1-score is obtained from the validation data. The important point here is that not only positive data (e.g., with cancerous regions), which can be used for normal training, but also negative (i.e., healthy, control) data, which are generally available in large quantities, can be used to detect hard boundary boxes, i.e., can be used for training.
@@
Figure 4: Positive example and five typical false positives provided by experts.
@@ 4 Experiments
4.1 Datasets
The image set used in this study consisted of 4,724 gastric X-ray images (1,117 images with 1,504 lesion annotations by radiologists and 3,607 images without lesions) from 145 patients in clinical settings provided by the Tokai University School of Medicine, Japan. A total of 116 images from five randomly selected patients (49 of them with 49 annotated boxes) were used as validation data to determine the hyperparameters and the decision to terminate HBBT, and the remaining 4,608 images from 140 patients were used to train and evaluate the proposed system.
Each image was in 8-bit grayscale and had a resolution of 2,048 × 2,048 pixels. Based on each expert annotation of the gastric cancer locations, the smallest rectangle surrounding each location was defined as a bounding box to be detected. A patient-based five-group cross-validation was used for training and evaluation. Precision, recall, and F1-score were used as the performance indicators of this system. The system provided candidate lesion bounding boxes and their associated confidence scores were above a predefined detection threshold $\xi$.
@@ 4.2 Details of the detection model
We used EfficientDet-D7 (Tan et al., 2020) as a gastric cancer detection model with an input resolution of 1,536 $\times$ 1,536 pixels. One difference from the original image size was the margin for random cropping performed during training. The entire networks are trained end-to-end and, as a result, can simultaneously estimate the detection target at any location and size in the input image and its content (i.e., ”cancer”). See the original paper for details on the model. The learning rate was 0.001, the batch size was 32, and the number of trainings was 100. The optimization method used was momentum SGD (Qian, 1999), and the exponential decay rate of the first-order moment was set to 0.9. To construct the proposed gastric cancer detection system, the proposed R-sGAIA was applied as an online augmentation to train this model, and healthy control images without cancer lesions, which could not be used in general object detection models, were also used for training by using HBBT.
@@ 4.3.2 Detection of suspicious cancerous regions
To confirm the effectiveness of the proposed system with R-sGAIA and HBBT, we conducted comparative experiments using the following combinations of conditions. Since there is no directly comparable literature on gastric cancer detection, we implemented Laddha et al.’s (2019) method based on YOLO-v3 and compared the performance on our task.
@@
    We selected two deep learning-based object detection models for this task: ResNet-101+Faster R-CNN (model 1) and EfficientDet-D7 (model 2). The first model has been widely used in the literature, including sGAIA. The second model is one of the advanced object detection models used in many studies, although newer models are currently being proposed. RandAugment (Cubuk et al., 2020), one of the state-of-the-art online augmentation methods, was applied to these models and used as the baseline for each model (i.e., baseline models 1 and 2, respectively). Here, color augmentation was omitted because this task deals with grayscale images. Based on the detection performance of the validation dataset, 4 of the 13 augment types were selected with an intensity level of 3 (i.e., $M=4$ and $N=3$ according to their notation).
@@
    We compared the performance of each model with and without the proposed technique. Regarding data augmentation, we chose either RandAugment, conventional sGAIA (+sGAIA), or the proposed R-sGAIA (+R-sGAIA). Regarding the proposed HBBT, this was annotated as use (+HBBT) or not.
@@ 5 Results
Figure 5 shows an example of the augmented image obtained with the proposed R-sGAIA for the leftmost original image. Figure 6 shows an example of the comparison of cancer detection on the EfficientDet-D7 network (baseline model 2) with and without the proposed R-sGAIA and HBBT. Here are the input images (a), the detection results with RandAugment (b), and those with the proposed R-sGAIA+HBBT (c). The proposed system (c) could accurately detect the gastric cancer region without being affected by the unevenness of the mucosa, and small areas as shown in the top and middle rows, respectively. In contrast, as shown in the bottom row, the result almost covers the ROI but is misaligned; therefore, we evaluated it as a false-positive.
The detection performance of gastric cancer regions from gastric X-ray images is summarized in Table 2. The object detection model detects regions where the probability (i.e., the confidence score) of belonging to the class of interest exceeds a predefined detection threshold of the bounding box $\xi$ as locations containing objects of that class. HBBT assigns new negative labels to the over-detected regions that exceed the threshold $\xi$ and retrains the model. Therefore, to draw an ROC curve for evaluation, the model must be trained every time $\xi$ is changed, which is expensive. In this experiment, therefore, the detection threshold $\xi$ was determined such that the SE (= recall) of detecting gastric cancer was about 90%, which was higher than the doctor’s 85.5%. At this detection threshold, HBBT improves precision by 2.3 points in Faster R-CNN (baseline model 1) and 3.3 points in EfficientDet-D7 (baseline model 2). Our system using the proposed techniques, R-sGAIA and HBBT, achieved a recall of 90.2%, a precision of 42.5%, and an F1-score of 0.578 (Table 2).
@@
The processing time required to process one X-ray image is approximately 0.51 s, thus making the proposed system a suitable practical screening system for gastric X-ray images.
Figure 6: Example of cancer detection on EfficientDet-D7 networks: (a) ground-truth region, (b) results obtained with RandAugment (baseline model 2), and (c) those with R-sGAIA+HBBT (proposed). The proposed R-sGAIA+HBBT led to the accurate detection for the top and middle examples, whereas for the bottom result, although it showed some improvement, the detection was inaccurate.
@@ 6 Discussion
The cancer detection performance of the proposed system in this study was significantly better than those achieved using the baseline methods, including RandAugment (a state-of-the-art general data augmentation strategy), the existing methods such as sGAIA Okamoto et al. (2019), and those proposed by (Laddha et al., 2019). In addition, its cancer detection capability was higher than that of the physician’s reading ($\sim$85.5%). In terms of percentage, more than 2 out of 5 detection results presented were cancer. From the standpoint of a physician using the system, this means that the system can provide a higher diagnostic yield than a conventional physician’s reading simply by checking only the areas surrounded by the few bounding boxes presented by the system in a gastric X-ray image. In other words, the system is practically effective for on-site medical examinations and can significantly reduce the burden on physicians.

### v4

@@ 2.3 Hard negative mining
Hard sample or hard example mining, a technique that improves model performance by explicitly retraining the model on error-prone or hard-to-identify data (hard samples), is widely used in machine learning (Felzenszwalb et al., 2009; Shrivastava et al., 2016). It is particularly effective in addressing class imbalance and has been extensively applied, alongside metric learning, in tasks such as face or person detection from complex backgrounds (Zhang et al., 2016; Smirnov et al., 2018; Chen et al., 2020), among other applications.
In recent years, deep learning-based object detection models have also adopted this strategy, reporting significant performance improvements by relearning misdetected bounding box regions, including in the medical field (Li et al., 2019; Tang et al., 2019). However, such object detection models typically cannot utilize images that do not contain any detection targets (such as background images or healthy control images) for training, since the error function to be minimized is defined based on the position and size of the bounding box for each detection target. Existing hard sample mining methods have been applied to data with at least one annotation box of the target class and have used overdetected or undetected regions or image patches as hard samples. Our HBBT is a form of hard sample mining or hard negative mining, which retrains regions of false positives. However, it differs significantly in that it allows the use of healthy control cases, which could not be used in traditional object detection models, for training. This is a particularly important contribution in the medical field, where the cost of acquiring positive samples is high.
@@ 3 Practical gastric cancer screening system
This study proposed a practical gastric cancer screening system using gastric X-ray images. The system highlights suspected cancerous sites with bounding boxes, allowing physicians to confirm the validity of the estimations. The two technical innovations in this study significantly improved system performance and achieved practical accuracy. The overall structure of the proposed system is shown in Figure 1. This study was approved by the Institutional Review Board of Tokai University Hospital (No. 20R-033, approval date: June 12, 2020). The main technical contributions were (1) R-sGAIA, which generates a diverse set of images that effectively emphasize the gastric fold region for data augmentation, and (2) HBBT, which reduces false positives by assigning “hard-sample” classes to bounding box regions where false positives are detected and iteratively learning from them. The object detection model used is EfficientDet (Tan et al., 2020), which incorporates EfficientNet (Tan and Le, 2019), a CNN known for its excellent performance, as its backbone, although the choice of model was arbitrary. The input images were 2,048 $\times$ 2,048 pixels, and the output consisted of a bounding box surrounding the presumed cancerous region in the image along with its confidence score.
@@ 3.1 R-sGAIA
The proposed R-sGAIA is a refined version of our previous method, sGAIA (Okamoto et al., 2019). Both R-sGAIA and sGAIA are probabilistic online data augmentation techniques designed to generate various images that highlight gastric folds in X-ray images, using medical knowledge to detect inflammation on the gastric mucosal surface. The original sGAIA improved gastric cancer detection by 6.9% in terms of F1-score (recall 92.3%, precision 32.4%) when applied to a Faster R-CNN model with ResNet-101 as the backbone, validating its effectiveness for practical screening. However, in sGAIA, the emphasis probability of each pixel $p(x,y)$, crucial for effective enhancement, was subjectively determined as a discrete value based on preliminary experiments. In the proposed R-sGAIA, $p(x,y)$ is instead modeled as a flexible continuous function, determined by only two hyperparameters. A schematic of the proposed R-sGAIA process is shown in Figure 2. Similar to sGAIA, R-sGAIA consists of four steps, with step 2 being the only difference.
@@
    $=$ $\frac{1}{1+\exp(-\gamma(E(x,y)-\theta))}.$
@@
Here, $\alpha$ and $\beta$ are augmentation parameters. The value of $\alpha$ ranges from 0.9 to 1.0, and $\beta$ ranges from --15 to --5, with both values chosen randomly for each processing. Based on this, the proposed R-sGAIA probabilistically enhances the location and intensity of the gastric fold region each time, generating a different enhanced gastric image for each trial. This makes it highly compatible with online data augmentation in deep neural network training. Another significant advantage of R-sGAIA is that it does not require any additional runtime. We applied this R-sGAIA for online augmentation in training a machine learning model (EfficientDet-D7) used for cancer region detection.
@@ 3.2 HBBT
The diagnostic support system for gastric X-ray images is a one-class detection problem focused solely on positive (i.e., malignant) regions. However, our task involves many regions with similar image features, making them difficult to distinguish, even for experts. With a large number of training images, recent deep learning-based object detection algorithms can achieve the desired detection performance for this task. However, this is challenging when labeled supervisory data is limited. Figure 4 shows typical false positives identified by experts in experiments using a vanilla detection model (i.e., EfficientDet-D7 without the proposed HBBT). The results indicate that most of these false positives fall into the five categories shown in examples (b)–(f) in Figure 4. However, it is difficult to clearly define and control these features in advance. In a fine-grained task like ours, it is crucial to reduce such hard negatives that exhibit high image similarity but are not the target. Note that our system detects a single class of candidate gastric cancer regions, with each detection box labeled as ”malignant.”
@@
One promising approach to address this issue is hard sample mining, specifically hard negative mining, which aims to reduce overdetection and has been extensively studied, as mentioned earlier. However, training hard negative mining in object detection models typically focuses on reducing false positives (hard negatives) outside the target region in the training image. The data available for training are constrained by the object detection model, meaning that only images with some object detection annotation label can be used, unless a framework like semi-supervised learning is employed.
@@
As mentioned earlier, a significant area for improvement in object detection models is the active use of negative samples, i.e., healthy controls. Conventional deep learning-based object detection algorithms typically only use images containing the object to be detected, excluding control cases without the object from training. This is a critical limitation. To address this, we propose HBBT, a simple yet effective method for suppressing false positives by incorporating the concept of hard negative mining. Training negative samples, which are more prone to false positives among the training data, is generally effective in constructing detectors with fewer false positives. In HBBT, as shown in Figure 1, the false-positive bounding boxes detected by the model are called hard boundary boxes. These are assigned hard-sample labels, converting the task into a two-class detection problem. The newly added hard-sample class label is treated as a negative example, and training is repeated until an optimal F1-score is achieved on the validation data. Importantly, this approach allows the use of not only positive data (e.g., images with cancerous regions) but also negative (i.e., healthy control) data, which are generally available in large quantities, for training by detecting and utilizing hard boundary boxes.
@@
Figure 4: Positive example and five typical false positives provided by experts (as shown in red box).
@@ 4 Experiments
4.1 Datasets and evaluation criteria
The image set used in this study consisted of 4,724 gastric X-ray images (1,117 images with 1,504 lesion annotations by radiologists and 3,607 images without lesions) from 145 patients in clinical settings, provided by the Tokai University School of Medicine, Japan. A total of 116 images from five randomly selected patients (49 of them with 49 annotated boxes) were used as validation data to determine the hyperparameters and to decide when to terminate HBBT. The remaining 4,608 images from 140 patients were used to train and evaluate the proposed system. Please note that we are not permitted to disclose the image data used in this experiment. Each image was an 8-bit grayscale with a resolution of 2,048 $\times$ 2,048 pixels. Based on expert annotations of the gastric cancer locations, the smallest rectangle surrounding each location was defined as a bounding box to be detected.
A subject-based five-fold cross-validation was used for training and evaluation, with precision, recall, and F1-score serving as the performance indicators for this system. The system provided candidate lesion bounding boxes, with their associated confidence scores exceeding a predefined detection threshold $\xi$. In the subject-based five-fold cross-validation, subjects were randomly divided into five groups. Images from four groups were used for training, while images from the remaining group were used for evaluation. This process was repeated five times, with a different group selected for evaluation each time. In medical research, including this experiment, data from the same subject can be inherently similar. Using standard cross-validation could lead to a virtual data leak, resulting in an overestimation of model performance due to the similarity between training and test data. To prevent this and ensure a rigorous evaluation, we used this evaluation strategy.
@@ 4.2 Details of the detection model
We used EfficientDet-D7 (Tan et al., 2020) as a gastric cancer detection model with an input resolution of 1,536 $\times$ 1,536 pixels. One difference from the original image size was the margin for random cropping performed during training. The entire network was trained end-to-end, allowing it to simultaneously estimate the detection target at any location and size within the input image and classify its content (i.e., ’cancer’). For details on the model, please refer to the original paper. The learning rate was set to 0.001, with a batch size of 32 and 100 training epochs. The optimization method used was momentum SGD (Qian, 1999), with the exponential decay rate of the first-order moment set to 0.9. To construct the proposed gastric cancer detection system, we applied the proposed R-sGAIA as an online augmentation technique during training. Additionally, healthy control images without cancer lesions, which are typically not used in general object detection models, were incorporated into the training process using HBBT.
@@ 4.3.2 Detection of suspicious cancerous regions
To confirm the effectiveness of the proposed system with R-sGAIA and HBBT, we conducted comparative experiments using the following combinations of conditions. Since there is no directly comparable literature on gastric cancer detection, we implemented Laddha et al.’s (2019) method based on YOLOv3 and compared the performance on our task.
@@
    We selected two deep learning-based object detection models for this task: ResNet-101+Faster R-CNN (model 1) and EfficientDet-D7 (model 2). The first model has been widely used in the literature, including in sGAIA studies. The second model is one of the advanced object detection models frequently employed in many studies, although newer models have since been proposed. The training images used in the experiment to detect candidate cancer regions were grayscale, not color images, and since there was no suitable pre-trained model available, we trained the model from scratch using He’s initial weights He et al. (2015) with the random seed set to ’2021’. RandAugment (Cubuk et al., 2020), one of the state-of-the-art online augmentation methods, was applied to these models and served as the baseline for each (i.e., baseline models 1 and 2, respectively). Color augmentation was omitted because this task involves grayscale images. Based on the detection performance of the validation dataset, 4 of the 13 augment types were selected with an intensity level of 3 (i.e., $M=4$ and $N=3$ according to their notation).
@@
    We compared the performance of each model with and without the proposed techniques. For data augmentation, we used either RandAugment, conventional sGAIA (+sGAIA), or the proposed R-sGAIA (+R-sGAIA). The use of the proposed HBBT was indicated as either applied (+HBBT) or not.
4.3.3 Computational environment for model training and evaluation
In all experiments, we used a computer equipped with a Xeon E5-2650v4 2.2GHz CPU, 256GB of memory, and a GeForce RTX 2080 Ti GPU, running on Ubuntu 20.04 LTS. PyTorch was used as the deep learning framework. The implementations of Faster R-CNN***https://pytorch.org/vision/master/models/faster_rcnn.html, EfficientDethttps://github.com/rwightman/efficientdet-pytorch, and RandAugmenthttps://pytorch.org/vision/main/generated/torchvision.transforms.RandAugment.html were sourced from the links provided. Please note that the OS versions listed here reflect those in use at the time of the initial submission and are not the latest versions.
@@ 5 Results
Figure 5 shows an example of the augmented image obtained with the proposed R-sGAIA for the leftmost original image. Figure 6 compares cancer detection on the EfficientDet-D7 network (baseline model 2) with and without the proposed R-sGAIA and HBBT. The input images are shown in (a), the detection results with RandAugment in (b), and those with the proposed R-sGAIA+HBBT in (c). The proposed system (c) accurately detected the gastric cancer region without being affected by mucosal irregularities or small areas, as demonstrated in the top and middle rows. However, as shown in the bottom row, the result nearly covers the ROI but is slightly misaligned, leading to a false-positive evaluation.
The detection performance of gastric cancer regions from gastric X-ray images is summarized in Table 2. The object detection model detects regions where the probability (i.e., the confidence score) of belonging to the class of interest exceeds a predefined detection threshold of the bounding box $\xi$ as locations containing objects of that class. HBBT assigns new negative labels to the over-detected regions that exceed the threshold $\xi$ and retrains the model. Therefore, to draw an ROC curve for evaluation, the model must be trained every time $\xi$ is changed, which is expensive. In this experiment, therefore, the detection threshold $\xi$ was determined such that the SE (= recall) of detecting gastric cancer was about 90%, which was higher than the doctor’s 85.5%. At this detection threshold, HBBT improves precision by 2.3 points in Faster R-CNN (baseline model 1) and 3.3 points in EfficientDet-D7 (baseline model 2). Our system, using the proposed techniques, R-sGAIA and HBBT, achieved a recall of 90.2%, a precision of 42.5%, and an F1-score of 0.578 (Table 2).
@@
The processing time required for one X-ray image is approximately 0.51 seconds, making the proposed system a practical and efficient screening tool for gastric X-ray images.
Figure 6: Example of cancer detection on EfficientDet-D7 networks (as shown in red box): (a) ground-truth region, (b) results obtained with RandAugment (baseline model 2), and (c) those with R-sGAIA+HBBT (proposed). The proposed R-sGAIA+HBBT led to the accurate detection for the top and middle examples, whereas for the bottom result, although it showed some improvement, the detection was inaccurate.
@@ 6 Discussion
The cancer detection performance of the proposed system in this study was significantly better than that achieved using baseline methods, including RandAugment (a state-of-the-art general data augmentation strategy), existing methods such as sGAIA Okamoto et al. (2019), and those proposed by (Laddha et al., 2019). Additionally, the system’s cancer detection capability exceeded that of a physician’s reading ($\sim$85.5%). In more than 2 out of 5 cases, the detection results presented by the system were cancerous. For a physician using the system, this means that by simply checking the areas surrounded by the few bounding boxes presented in a gastric X-ray image, a higher diagnostic yield can be achieved than with a conventional reading. In other words, the system is practically effective for on-site medical examinations and can significantly reduce the burden on physicians.